OpenStudy (anonymous):

write an equation of the line through (-4,3) that is perpendicular to the line 2x-y+3=0

OpenStudy (anonymous):

so if the slope is 2 then the perpendicular slope is -1/2

OpenStudy (amistre64):

2x-y+3=0 ; get rid of the +3 and 0 2x-y=n now switch x and y coeffs -x+2y = n ; and change out the operator -x-2y = n ; fill in your point to calibrate the equation and your done

OpenStudy (anonymous):

y-3=-1/2(x+4) -2y+6=x+4 -2y-x=-2

OpenStudy (amistre64):

-x-2y = n ; (-4,3) -(-4)-2(3) = n 4 -6 = n = -2 -x-2y = -2 -x-2y +2 = 0

OpenStudy (anonymous):

same thing yay! thanks amistre

OpenStudy (amistre64):

since slope is defined by your x and y coordinates; just flip them and negate one :)

OpenStudy (anonymous):

lol okay